Given the following cash inflow at the end of each​ year, what is the future value of this cash flow at 3 %​, 8 %​, and 16​% interest rates at the end of year​ 7?

Year​ 1: ​$13 comma 000

Year​ 2: ​$22 comma 000 Year​

3: ​$32 comma 000 Years

4 through​ 6: ​$0 Year​

7: ​$140 comma 000

What is the future value of this cash flow at 3​% interest rate at the end of year​ 7?
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

At 3%:

A=13000*(1.03)^6+22000*(1.03)^5+32000*(1.03)^4+140,000

=$217042.99(Approx).

At 8%:

A=13000*(1.08)^6+22000*(1.08)^5+32000*(1.08)^4+140,000

=$236,490.23(Approx).

At 16%:

A=13000*(1.16)^6+22000*(1.16)^5+32000*(1.16)^4+140,000

=$275,821.13(Approx).